OpenStudy (anonymous):

The ratio of the radii of two spheres is 3:1. The volume of the larger sphere is 729 in3. What is the volume of the smaller sphere

OpenStudy (anonymous):

i believe the equation for the volume of a sphere is \[v = \pi r^3\] What you can do is plug 729 into that equation for v, solve for r and divide the answer by 3. Then, plug the answer you get from that into the Volume equation and you'll get the volume of the smaller sphere

OpenStudy (anonymous):

Did this help?

OpenStudy (ipwnbunnies):

The equation is actually \[V(sphere) = \frac{ 4 }{ 3 }*\pi*r^{3}\]

OpenStudy (anonymous):

yeah, listen to him, i'm sure he knows more of this than i do.

OpenStudy (radar):

If you do the math you will see the answer is a whole number (2 digits).

OpenStudy (radar):

Since the volume formula would only be different in radius and since the radii is cubed you could take the given ratio (3:1) and cube it 27:1. The smaller sphere then would have as you can see have 1/27 of the volume of the larger sphere. That method keeps Pi out of it. lol

OpenStudy (radar):

Sorry about the deleted post where I squared the ratio rather that cube it. Now you only need to divide 729 in^3 by 27 to get the number of cubic in. in the smaller sphere.